# A Survey of Low-bit Large Language Models: Basics, Systems, and Algorithms

### v2

@@ 3.1.1 Ready-to-use Algorithms
With the emergence of quantization algorithms for LLMs, some typical methods have already been integrated into most frameworks, while some methods may be developed and published originally on a specific framework. We list the most ready-to-use algorithms in each mainstream framework in Table 2. Some methods are included by most frameworks, such as GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2024a), SmoothQuant (Xiao et al., 2023), and so on. These methods share several advantages: high accuracy and efficient performance after quantization, seamless integration into existing implementation procedures, and user-friendliness.
In addition, some algorithms are supported by several frameworks. For example, LLM.int8() (Dettmers et al., 2022b) was well supported by bitsandbytes (in HuggingFace), which allows to store and load 8-bit weights directly from the HuggingFace Hub and quantize weight in linear layers to 8-bit. FP6-LLM (Xia et al., 2024) is integrated in DeepSpeed-FastGen [^24] https://github.com/microsoft/DeepSpeed/tree/master/blogs/deepspeed-fastgen (Holmes et al., 2024) to implement the runtime quantization for 6-bit floating-point weight-only quantization. It allows efficient quantization and dequantization of 6-bit weight LLMs through a unified configuration option. It is noteworthy that Transformers (by HuggingFace) and QServe (by MIT EECS Lin et al. (2024b)) integrate most algorithms with comprehensive user manuals and detailed examples, enabling a quick start for deep learning researchers and developers.
@@ 3.1.2 Bitwidth Support
Weight-onlybit means only quantizing the weight while keeping FP16 activation (Lin et al., 2024a). The quantized weight will be dequantized back to FP16 using pre-obtained scaling factors, and then conduct FP16 mma with FP16 activation. Therefore, it theoretically supports non-uniform quantization with arbitrary bitwidth. The speedup is achieved by reducing the latency of data transmission between the computing device and storage host with smaller amounts of weight data, but the dequantizing of weight costs extra time. The detailed speedup will be discussed in Section 3.2.1.
Wbit&Abit means that the algorithm quantizes both the weight and activation, and conducts low-bit matrix multiplication (MatMul) in low-level (for example, in PTX ISA 8.5252525https://docs.nvidia.com/cuda/parallel-thread-execution/index.html for NVIDIA GPUs, instruction mma.sync.aligned.shape.row.col .s32.u4.u4.s32 means the data type of the multipliers is the 4-bit unsigned integer). All the frameworks support the INT8 and FP16 MatMul. However, limited by the computing capabilities of the hardware and the supported operations in the instruction set, only part of them have INT4 and FP8 MatMul. Few supports different bitwidth of weight and activation (like $W_{\mathrm{int4}}A_{\mathrm{int8}}$), which requires customized computation kernels with assembled GEMV instructions [^26] https://huggingface.co/docs/transformers/main/en/quantization/eetq  (Egiazarian et al., 2024a). It should be mentioned that if you want to use low-bit MatMul, your hardware architecture must support the specific low-bit computing, and it is necessary to upgrade/downgrade the driver to the corresponding version to reproduce the real low-bit computation and get the desired speedup ratio.
KV Cachebit lists the bitwidth of Key-Value Cache. As a caching technology, memory consumption of the KV cache increases rapidly as batch size and sequence length continue to grow, potentially surpassing the model size. Therefore, quantizing the KV cache significantly reduces memory usage during model inference. There are several works devoted to quantizing the KV cache (Hooper et al., 2024; Yue et al., 2024; Liu et al., 2024c). Similar to weight-only algorithms, the quantized key-value pairs usually need to be dequantized to floating-point before MatMul, otherwise, the specific system support of multiplying low-bit to floating-point is required. Except for the listed bitwidth, all frameworks support the FP16 KV cache, which means directly storing the activation.
@@ 3.2 System Support for Quantization
In practical implementations, it is perplexing that some quantization algorithms, although reducing the bitwidth of weight or activation, do not lead to faster inference. Therefore, a critical question comes into mind: How do quantization actually achieve real acceleration and storage saving? To answer this question, we first need to clarify the data transmission process involved in model inference.
Figure 4: Data transmission of weight and activation in the caching system during inference. The bandwidth and latency are officially reported by NVIDIA A100 as an example.
@@
Host memory $\rightarrow$ Device memory. For weight, we load one layer’s weight from the host memory to the device’s global memory. The bandwidth is relatively low, which is 25 GB/s per direction (taking NVIDIA A100 as an example (Smith, 2020)). If quantized, it is always in a compact format, thus the time can be saved. The activation is originally generated on the device during inference, which does not need to be copied from the host.
@@ 3.2.1 Weight-Only Quantization
The fundamental bottleneck in model inference before and after the advent of large models is the data transmission and storage costs, which are always neglected in ordinary small models. Due to the large amount of data, the transmission latency can not be overlooked, which even surpasses the computation latency and becomes the major challenge in LLM inference. Therefore, weight-only quantization emerges, which compacts the weight and reduces the data copy burden among levels of caches (Lin et al., 2024a; Frantar et al., 2022).
@@ 3.2.3 KV Cache Quantization
There are mainly three techniques for quantizing KV cache: (1) Quantizing to lower bitwidth. QoQ (Lin et al., 2024b) compresses KV to 4-bit and proposes SmoothAttention to prevent the accuracy drop due to the lower bitwidth. KIVI (Zirui Liu et al., 2023) even developed a tuning-free 2bit KV cache quantization algorithm. Yang et al. (2024) proposes a mixed-precision strategy that quantizes the earliest KV to lower bitwidth, while keeping the new KV with more bits. (2) Quantizing window. Many studies postpone the quantization of KV pairs, but only quantize them in a batch when the length of the full-precision KV list exceeds the window size. (3) Skipping the dequantization of $\mathbf{K}_{new}$. Methods like WKVQuant (Yue et al., 2024) concatenates the FP $\mathbf{K}_{new}$ and $\mathbf{V}_{new}$ to the dequantized $\mathbf{K}_{prev}$ and $\mathbf{V}_{prev}$, which preserves more information of current token in $\mathbf{K}$ and $\mathbf{V}$ matrices, then quantizes the $\mathbf{K}_{new}$ and $\mathbf{V}_{new}$ and stores them into the KV cache (when meets the condition). (4) Optimizing outliers. There are token-wise outliers in KV matrices, so methods such as storing the outliers with higher bitwidth or mitigating the outlier magnitudes can improve the performance (Dong et al., 2024; Liu et al., 2024c; Kang et al., 2024; Lin et al., 2024b). We omit the details of this category, as the general practices are similar to the quantization methods used for the entire model.
@@ 3.2.4 Quantization and Dequantization
Quantizing higher bitwidth floating-point to lower is actually the clip of mantissa bits. That is because the source value with higher bitwidth usually has more or equal bits for both exponent and mantissa parts compared to the target value with lower bitwidth. Algorithm 1 provides an example of quantizing FP32 to FP8. And we follow Micikevicius et al. (2022) to summarize the general process as follows:
@@
[…]
@@
2:  $e^{min}=$(1 << ($e$ 1)) + 1
3:  $e^{max}=$ (1 << ($e$ 1))
4:  $m=x-e-1$ // Theoretical maximum of exponent part for FP8
@@
6:  ${X_{\mathrm{FP}8}^{m}}=$(0x007FFFFF >> $m$) & 0x007FFFFF
@@
13:     $X_{\mathrm{FP}8}^{sign}=X_{\mathrm{FP32}}^{unscaled}$ 0x80000000
14:     $X_{\mathrm{FP}8}^{e}=X_{\mathrm{FP32}}^{unscaled}$ 0x7F800000
15:     $X_{\mathrm{FP}8}^{m}=X_{\mathrm{FP32}}^{unscaled}$ 0x007FFFFF // Check exponent underflow
@@
20:     $R_{m}=(X_{\mathrm{FP}8}^{m}\texttt{<<}m)$ 0x007FFFFF + 0x3F800000
@@
We first scale the floating-point numbers to the representation span of $\mathrm{INT}k$ by dividing the scaling factor $s\in\mathbb{R^{+}}$, and adding a zero-point $z\in\mathbb{Z}$ to shift the clamped range (Wu et al., 2020). $\mathrm{round}(\cdot)$ is the round-to-the-nearest function, and $\mathrm{clamp}(\cdot,q^{\min},q^{\max})$ restricts values to be within the representation span of $k$-bit with $q^{\min}=-2^{k-1},q^{\max}=2^{k-1}-1$ in symmetric quantization and $q^{\min}=0,q^{\max}=2^{k}-1$ in asymmetric quantization. Therefore, the overall quantization formulation can be written as:
@@
where the scaling factor $s$ can be initialized as $s_{0}=({q^{\max}-q^{\min}})/({X_{\mathrm{FP}}^{\max}-X_{\mathrm{FP}}^{\min}}),$ where $X_{\mathrm{FP}}^{\max}$ and $X_{\mathrm{FP}}^{\min}$ are the maximum and minimum values.
For system support, many frameworks apply the Marlin quantization [^29] https://github.com/IST-DASLab/marlin (Frantar and Alistarh, 2024) as the standard process. The pseudocode Algorithm 2 outlines the steps involved in Marlin quantization, and uses 4-bit integer quantization as an example. The values are quantized and stored as unsigned integers with the desired bitwidth. Extra pre/post-shift will be conducted to get the signed values. Therefore, we first scale the $X_{\mathrm{FP32}}$ values by $s$ and round it to integers. Then, adding $2^{k-1}$ to shift the values to non-negative integers within the span of uINT4 (4-bit unsigned integer). We omit the detail of C++ built-in datatype casting function float2uint. To be understood, we explain the packing process in lines 4 to 8 by double for loops, which is actually implemented as lines 9 to 11. In 4-bit quantization, every 8 values are packed as a single uINT32, and the quantized matrix size is a quarter of the original. By using $i$::8, we abstract every 8 values along the dimension $C$ starting with $i$, incrementing by 8, and ending by default (till the end of dimension $C$). And then left shift the values by $4*i$ to place the 4-bit value to the corresponding bit range, and leave $4*i$ zeros on the right, allowing previously-stored quantized values to be preserved after OR operation.
@@
8:  end for// Line 4$\sim$8 can also simplify as line 9$\sim$11 // i::8 creates a sequence starts at $i$, increments by $8$, and ends by default
9:  for $i\leftarrow 0$ to $7$ do
@@
There are several custom algorithms that introduce faster data type conversions. QQQ (Zhang et al., 2024d) designs a faster FP16 to INT8 conversion, named FastFP16toINT8. It starts by shifting the FP16 values to the representation span of uINT8 by adding 128. Next, adding an additional 1024 which effectively converts and places the 8 bits of uINT8 into the lower segment of the FP16 mantissa. Finally, the lower 8 bits from FP16 are extracted and applied with an XOR operation with 0x80 to obtain the desired INT8 format. The overall process can be further simplified to FMA, PRMT, and XOR operations in practice.
@@
Therefore, in many works $s$ can also be initialized by searching from candidates to find an optimal (Wei et al., 2023b):
$s_{\mathrm{candidate}}$ $=\frac{i}{num_{i}}s_{0},\quad i\in\mathbb{Z^{+}},i\in(0,num_{i})$ (10)
$s.t.~{}\min||{X}_{\mathrm{FP}}-\hat{X}_{\mathrm{FP}}||_{p}.$ (11)
where $num_{i}$ means the number of candidates, which is always set as 50, 100 and so on (Yuan et al., 2024; Wei et al., 2023b). $s$ can also be a learnable parameter (Wei et al., 2023b; Shao et al., 2023). The way to find a better $s$ has been widely studied before LLMs emerged (Ding et al., 2024; Wei et al., 2023a).
For system support, we first unpack the elements according to the way we pack them, and then multiply them to the corresponding scaling factor, which can be tensor-wise, channel-wise, token-wise, and other granularities described in Section 2.2. Custom implementations are also proposed, SINT4toS8 Li et al. (2023a) designs a faster conversion from INT4 to INT8 by multiplying by 16.
@@
Using sign or bool depends on the algorithm design, i.e. what value we expect the bits to represent. For example, binarized transformers always use bool function on attention scores and the post-ReLU activation. While the weight and activation in linear functions take sign function. Since the hardware always regards the bits as 0 or 1, we can assemble instructions to achieve any desired matrix multiplication [^30] https://github.com/yifu-ding/BGEMM-CUDA. For example, on NVIDIA GPUs, the mma instruction takes 0/1 bit matrices and regards them as 0s and 1s while conducting bitwise accumulation operation popcount. Therefore, to obtain the correct accumulation, the popcount function is designed with different arithmetic rules, i.e., if it substracts 1 for each 0, we can have the result of sign function. It has many accelerated implementations, such as lookup table [^31] https://github.com/WojciechMula, nifty popcnt (Wilkes et al., 1958), hacker popcnt (Warren, 2012), hakmem popcnt  [^32] https://en.wikipedia.org/w/index.php?title=HAKMEM&oldid=1228234783 and so on.
@@
It is simply by multiplying a scaling factor $s$, i.e. $\hat{X}_{\mathrm{FP}}=s\cdot X_{\texttt{sign}/\texttt{bool}}$ to preserve the magnitude of the original values. It is easy to understand that large amounts of information will be lost in binarization. Therefore, few studies are devoted to binarizing LLMs due to the sharp performance degradation. Due to the significant speedup and storage reduction, it is valuable to dig deeper into binarizing LLMs, but may require a new formulation beyond sign and bool functions. DB-LLM (Chen et al., 2024a) proposes 2-bit weight quantization by decomposing to two 1-bit weight matrices, which can be efficient in MatMul theoretically.

### v3

@@ 3.1.1 Ready-to-use Algorithms
With the emergence of quantization algorithms for LLMs, some typical methods have already been integrated into most frameworks, while some methods may be developed and published originally on a specific framework. We list the most ready-to-use algorithms in each mainstream framework in Table 2. Some methods are included by most frameworks, such as GPTQ (frantar2022gptq), AWQ (lin2024awq), SmoothQuant (xiao2023smoothquant), and so on. These methods share several advantages: high accuracy and efficient performance after quantization, seamless integration into existing implementation procedures, and user-friendliness.
In addition, some algorithms are supported by several frameworks. For example, LLM.int8() (dettmers2022llmint8) was well supported by bitsandbytes (in HuggingFace), which allows to store and load 8-bit weights directly from the HuggingFace Hub and quantize weight in linear layers to 8-bit. FP6-LLM (xia2024fp6) is integrated in DeepSpeed-FastGen [^24] https://github.com/microsoft/DeepSpeed/tree/master/blogs/deepspeed-fastgen (holmes2024deepspeedfastgenhighthroughputtextgeneration) to implement the runtime quantization for 6-bit floating-point weight-only quantization. It allows efficient quantization and dequantization of 6-bit weight LLMs through a unified configuration option. It is noteworthy that Transformers (by HuggingFace) and QServe (by MIT EECS lin2024qserve) integrate most algorithms with comprehensive user manuals and detailed examples, enabling a quick start for deep learning researchers and developers.
@@ 3.1.2 Bitwidth Support
Weight-onlybit means only quantizing the weight while keeping FP16 activation (lin2024awq). The quantized weight will be dequantized back to FP16 using pre-obtained scaling factors, and then conduct FP16 mma with FP16 activation. Therefore, it theoretically supports non-uniform quantization with arbitrary bitwidth. The speedup is achieved by reducing the latency of data transmission between the computing device and storage host with smaller amounts of weight data, but the dequantizing of weight costs extra time. The detailed speedup will be discussed in Section 3.2.1.
Wbit&Abit means that the algorithm quantizes both the weight and activation, and conducts low-bit matrix multiplication (MatMul) in low-level (for example, in PTX ISA 8.5252525https://docs.nvidia.com/cuda/parallel-thread-execution/index.html for NVIDIA GPUs, instruction mma.sync.aligned.shape.row.col .s32.u4.u4.s32 means the data type of the multipliers is the 4-bit unsigned integer). All the frameworks support the INT8 and FP16 MatMul. However, limited by the computing capabilities of the hardware and the supported operations in the instruction set, only part of them have INT4 and FP8 MatMul. Few supports different bitwidth of weight and activation (like $W_{\mathrm{INT4}}A_{\mathrm{INT8}}$), which requires customized computation kernels with assembled GEMV instructions [^26] https://huggingface.co/docs/transformers/main/en/quantization/eetq  (egiazarian2024extremecompressionlargelanguage). It should be mentioned that if you want to use low-bit MatMul, your hardware architecture must support the specific low-bit computing, and it is necessary to upgrade/downgrade the driver to the corresponding version to reproduce the real low-bit computation and get the desired speedup ratio.
KV Cachebit lists the bitwidth of Key-Value Cache. As a caching technology, memory consumption of the KV cache increases rapidly as batch size and sequence length continue to grow, potentially surpassing the model size. Therefore, quantizing the KV cache significantly reduces memory usage during model inference. There are several works devoted to quantizing the KV cache (hooper2024kvquant; yue2024wkvquant; liu2024intactkv). Similar to weight-only algorithms, the quantized key-value pairs usually need to be dequantized to floating-point before MatMul, otherwise, the specific system support of multiplying low-bit to floating-point is required. Except for the listed bitwidth, all frameworks support the FP16 KV cache, which means directly storing the activation.
@@ 3.2 System Support for Quantization
In practical implementations, it is perplexing that some quantization algorithms, although reducing the bitwidth of weight or activation, do not lead to a faster inference. Therefore, a critical question comes into mind: How does quantization actually achieve real acceleration and storage saving? To answer this question, we must first clarify the data transmission process involved in model inference.
Figure 4: Data transmission of weight and activation in the caching system during inference. The bandwidth and latency are officially reported by NVIDIA A100 as an example. PCIe is a high-speed interface standard used for connecting various hardware components, such as GPUs, SSDs. Async_Copy means asynchronous data copy using cp.async intrinsic. ldmatrix and lds are data loading instructions that load matrix from shared memory to registers with a strict layout requirement or in a fine-grained and flexible manner, respectively (NVIDIA_PTX_ISA).
@@
Host memory $\rightarrow$ Device memory. For weight, we load one layer’s weight from the host memory to the device’s global memory. The bandwidth is relatively low, which is 25 GB/s per direction (taking NVIDIA A100 as an example (smith2020nvidia)). If quantized, it is always in a compact format, thus the time can be saved. The activation is originally generated on the device during inference, which does not need to be copied from the host.
@@ 3.2.1 Weight-Only Quantization
The fundamental bottleneck in model inference before and after the advent of large models is the data transmission and storage costs, which are always neglected in ordinary small models. Due to the large amount of data, the transmission latency can not be overlooked, which even surpasses the computation latency and becomes the major challenge in LLM inference. Therefore, weight-only quantization emerges, which compacts the weight and reduces the data copy burden among levels of caches (lin2024awq; frantar2022gptq).
@@ 3.2.3 KV Cache Quantization
There are mainly four techniques for quantizing KV cache: (1) Quantizing to lower bitwidth. QoQ (lin2024qserve) compresses KV to 4-bit and proposes SmoothAttention to prevent the accuracy drop due to the lower bitwidth. KIVI (kivi2023plug) even developed a tuning-free 2-bit KV cache quantization algorithm. yang2024no proposes a mixed-precision strategy that quantizes the earliest KV to lower bitwidth, while keeping the new KV with more bits. (2) Quantizing window. Many studies (zhang2024kv; duanmu2024skvq) postpone the quantization of KV pairs, but only quantize them in a batch when the length of the full-precision KV list exceeds the window size. For example, SKVQ (duanmu2024skvq) employs a sliding-window mechanism, determining the quantization parameters within the window. (3) Skipping the dequantization of $\mathbf{K}_{new}$. Methods like WKVQuant (yue2024wkvquant) concatenates the FP $\mathbf{K}_{new}$ and $\mathbf{V}_{new}$ to the dequantized $\mathbf{K}_{prev}$ and $\mathbf{V}_{prev}$, which preserves more information of current token in $\mathbf{K}$ and $\mathbf{V}$ matrices, then quantizes the $\mathbf{K}_{new}$ and $\mathbf{V}_{new}$ and stores them into the KV cache (when meets the condition). (4) Optimizing outliers. There are token-wise outliers in KV matrices, so methods such as storing the outliers with higher bitwidth or mitigating the outlier magnitudes can improve the performance (dong2024qaqqualityadaptivequantization; liu2024intactkv; kang2024gear; lin2024qserve). We omit the details of this category, as the general practices are similar to the quantization methods used for the entire model.
@@ 3.2.4 Quantization and Dequantization
Quantizing higher bitwidth floating-point to lower is actually the clip of mantissa bits. That is because the source value with higher bitwidth usually has more or equal bits for both exponent and mantissa parts compared to the target value with lower bitwidth. Algorithm 1 provides an example of quantizing FP32 to FP8. And we follow micikevicius2022fp8 to summarize the general process as follows:
@@
(2) Check Overflow/Underflow. Check whether the source value overflows the FP8 range, either from the upper or lower bound. If so, return the maximum or minimum directly. If it is not overflow, check if the exponent part underflows from the smallest positive normal number that the FP8 format can present. If so, we divide the value by the smallest subnormal number in FPxx, round to the nearest integer, and then multiply the smallest subnormal number. The integer determines the value of mantissa bits and the exponent bits are all set to zero.
@@
2:  $e^{min}=$ -(1 << ($e$ - 1)) + 1
3:  $e^{max}=$ (1 << ($e$ - 1))
4:  m=xe1m=x-e-1 // Theoretical maximum of exponent part for FP8
@@
6:  $X_{\mathrm{FP}8}^{m}$}= ~(0x007FFFFF >> $m$) & 0x007FFFFF
@@
13:   $X_{\mathrm{FP}8}^{sign}=X_{\mathrm{FP32}}^{unscaled}$ & 0x80000000
14:   $X_{\mathrm{FP}8}^{e}=X_{\mathrm{FP32}}^{unscaled}$ & 0x7F800000
15:   $X_{\mathrm{FP}8}^{m}=X_{\mathrm{FP32}}^{unscaled}$ & 0x007FFFFF // Check exponent underflow
@@
20:   $R_{m}=(X_{\mathrm{FP}8}^{m}\texttt{<<}m)$ & 0x007FFFFF + 0x3F800000
@@
We first scale the floating-point numbers to the representation span of $\mathrm{INT}k$ by dividing the scaling factor $s\in\mathbb{R^{+}}$, and adding a zero-point $z\in\mathbb{Z}$ to shift the clamped range (wu2020integerquantizationdeeplearning). $\mathrm{round}(\cdot)$ is the round-to-the-nearest function, and $\mathrm{clamp}(\cdot,q^{\min},q^{\max})$ restricts values to be within the representation span of $k$-bit with $q^{\min}=-2^{k-1},q^{\max}=2^{k-1}-1$ in symmetric quantization and $q^{\min}=0,q^{\max}=2^{k}-1$ in asymmetric quantization. Therefore, the overall quantization formulation can be written as:
@@
where the scaling factor $s$ can be initialized as $s_{0}=({X_{\mathrm{FP}}^{\max}-X_{\mathrm{FP}}^{\min}})/$ $({q^{\max}-q^{\min}}),$ where $X_{\mathrm{FP}}^{\max}$ and $X_{\mathrm{FP}}^{\min}$ are the maximum and minimum values.
For system support, many frameworks apply the Marlin quantization [^29] https://github.com/IST-DASLab/marlin (frantar2024marlin) as the standard process. The pseudocode Algorithm 2 outlines the steps involved in Marlin quantization, and uses 4-bit integer quantization as an example. The values are quantized and stored as unsigned integers with the desired bitwidth. Extra pre/post-shift will be conducted to get the signed values. Therefore, we first scale the $X_{\mathrm{FP32}}$ values by $s$ and round it to integers. Then, adding $2^{k-1}$ to shift the values to non-negative integers within the span of uINT4 (4-bit unsigned integer). We omit the detail of C++ built-in datatype casting function float2uint. To be understood, we explain the packing process in lines 4 to 8 by double for loops, which is actually implemented as lines 9 to 11. In 4-bit quantization, every 8 values are packed as a single uINT32, and the quantized matrix size is a quarter of the original. By using ii::8, we abstract every 8 values along the dimension $C$ starting with $i$, incrementing by 8, and ending by default (till the end of dimension $C$). And then left shift the values by 4i4*i to place the 4-bit value to the corresponding bit range, and leave 4i4*i zeros on the right, allowing previously-stored quantized values to be preserved after OR operation.
@@
[…]
9:  for $i\leftarrow 0$ to 77 do
@@
There are several custom algorithms that introduce faster data type conversions. QQQ (qqq) designs a faster FP16 to INT8 conversion, named FastFP16toINT8. It starts by shifting the FP16 values to the representation span of uINT8 by adding 128. Next, adding an additional 1024 which effectively converts and places the 8 bits of uINT8 into the lower segment of the FP16 mantissa. Finally, the lower 8 bits from FP16 are extracted and applied with an XOR operation with 0x80 to obtain the desired INT8 format. The overall process can be further simplified to FMA, PRMT, and XOR operations in practice.
@@
Therefore, in many works $s$ can also be initialized by searching from candidates to find an optimal (wei2023outlier):
$s_{\mathrm{candidate}}$ $=\frac{i}{num_{i}}s_{0},\qquad i\in\mathbb{Z^{+}},i\in(0,num_{i})$ (12)
[…]
where $num_{i}$ means the number of candidates, which is always set as 50, 100 and so on (ptq4vit; wei2023outlier). $s$ can also be a learnable parameter (wei2023outlier; shao2023omniquant). The way to find a better $s$ has been widely studied before LLMs emerged (ding2024reg; qdrop; tian2024qvd).
For system support, we first unpack the elements according to the way we pack them, and then multiply them to the corresponding scaling factor, which can be tensor-wise, channel-wise, token-wise, and other granularities described in Section 2.2. Custom implementations are also proposed, SINT4toS8 li2023speedodysseydeployablequantization designs a faster conversion from INT4 to INT8 by multiplying by 16.
@@
Using sign or bool depends on the algorithm design, i.e. what value we expect the bits to represent. For example, binarized transformers always use bool function on attention scores and the post-ReLU activation. While the weight and activation in linear functions take sign function. Since the hardware always regards the bits as 0 or 1, we can assemble instructions to achieve any desired matrix multiplication [^30] https://github.com/yifu-ding/BGEMM-CUDA. For example, on NVIDIA GPUs, the mma instruction takes 0/1 bit matrices and regards them as 0s and 1s while conducting bitwise accumulation operation popcount. Therefore, to obtain the correct accumulation, the popcount function is designed with different arithmetic rules, i.e., if it substracts 1 for each 0, we can have the result of sign function. It has many accelerated implementations, such as lookup table [^31] https://github.com/WojciechMula, nifty popcnt (wilkes1958preparation), hacker popcnt (warren2012hacker), hakmem popcnt  [^32] https://en.wikipedia.org/w/index.php?title=HAKMEM&oldid=1228234783 and so on.
@@
It is simply by multiplying a scaling factor $s$, i.e. $\hat{X}_{\mathrm{FP}}=s\cdot X_{\texttt{sign}/\texttt{bool}}$ to preserve the magnitude of the original values. It is easy to understand that large amounts of information will be lost in binarization. Therefore, few studies are devoted to binarizing LLMs due to the sharp performance degradation. Due to the significant speedup and storage reduction, it is valuable to dig deeper into binarizing LLMs, but may require a new formulation beyond sign and bool functions. DB-LLM (dbllm) proposes 2-bit weight quantization by decomposing to two 1-bit weight matrices, which can be efficient in MatMul theoretically.